6. Applying the rule of 3, determine the accuracy of the following specimens and state whether further action is indicated: 6pts. Case A: Hgb: 12g/dL Measured HCT: 35% Calculation: Further Action? Yes or No (circle one) Case B: Hgb: 9 g/dL Measured HCT: 32% Calculation: Further Action? Yes or No
Answer #1

Rule of three is used to check if the blood results of patient are valid or not. This rule is valid for only those specimens that have normocytic, normochromic erythrocytes. As per the rule the value of HCT (hematocrit) should be three times to the value of Hgb ( hemoglobin value), plus minus three.

Rule of Three: Hemoglobin x 3 = Hematocrit +/-3

Case A) Hgb =12g/dl

HCT= 35% or 0.35 L/L

Calculation = Hgb (12) x 3 = HCT (36)

The acceptable range for the HCT would be 33% to 39%. So actual hematocrit value (35%) is within the acceptable range. So no further action is required.

Case B)  Hgb = 9g/dl

HCT= 32% or 0.32 L/L

Calculation = Hgb (9) x 3 = HCT (27)

The acceptable range for the HCT would be 24% to 30%. Actual hematocrit value (32%) does not fall within the acceptable  range. So, yes further action is required